Suppose you sold a call option with a $50 premium. The strike price is $1,200. What is your expected payoff if the price of the underlying asset is $1,000.
Answer #1

A call option is exercised when the stock price is more than strike price. In this case the option will not be exercised as the stock price(1,000) is less than strike price (1,200).

Option writer payoff / profit

= Max ( 0, stock price - strike price) + call premium

= Max (0, 1,000 - 1,200) + 50

= Max (0, -200) + 50

= 0 + 50

= $50